OpenStudy (anonymous):

Thanks for helping! Complex conjugate of z2=8(cos(7pi/6)+isin(7pi/6)))

OpenStudy (anonymous):

@Hero @Michele_Laino @ganeshie8 @mathstudent55 @Nnesha @Loser66 @Luigi0210 @whpalmer4

OpenStudy (anonymous):

My worksheet says /z subscript 2. Which I think means complex conjugate.

OpenStudy (anonymous):

It gave me the z2 equation and just says /z2

OpenStudy (anonymous):

okay, thanks for attempting to help :)

OpenStudy (anonymous):

What does the x' and y' mean?

OpenStudy (anonymous):

oh, gotcha

OpenStudy (anonymous):

So the conjugate is 2 sqrt(7)?

OpenStudy (anonymous):

Okay, thanks! That was pretty helpful.

OpenStudy (irishboy123):

*if* \( z_2=8\ (cos\frac {7 \pi}{6} \ + \ isin\frac{7\pi}{6})\), then \( \bar z_2=8\ (cos\frac {7 \pi}{6} \ - \ isin\frac{7\pi}{6})\)

OpenStudy (anonymous):

Oh, so it only needs a subtraction?

OpenStudy (irishboy123):

@NotSephiroth yes \(z = x + i y\), \(\bar z = x - i y\)

OpenStudy (irishboy123):

or \(z = R e^{i \theta}\), \(\bar z = R e^{-i \theta}\)

OpenStudy (anonymous):

Okay thanks!

OpenStudy (anonymous):

that was really helpful!